TAU operational status:
Telemetry is down since day 2022,106 (4/16) due to issue with radio link. Station operator has been requested to investigate.
The station is back up after the station operator replaced the Newmar power supply shipped from UCSD.
The station is down as of day 2018,348 due to failure of the Newmar power supply.
The station operator visited the station and replaced the tau00 digitizer, and GPS antenna, restoring tau00 channels at 2018,277-03:08.
The station operator visited the station to replace the tau00 digitizer, and reports the GPS equipment is damaged. An IDA engineer is working with the station operator to resolve the GPS issue. The tau00 channels are not being acquired (as of day 2018,255-01:17) until the GPS issue is resolved.
IDA station TAU is experiencing problems with the GPS clocks. A new digitizer, GPS antenna, and cable are being prepared to ship to the station.
IDA station TAU is back online as of day 2015,024-03:35 after IDA engineers visited the station and upgraded the power system, installed a new auxiliary sensor, and re-installed the primary sensors adding thermal insulation.
IDA station TAU is down as of day 006-14:51 due to damage to the microwave link between the digitizers and the station computer. A trip by IDA engineers is planned for this month.
An IDA engineer visited the station to investigate telemetry issues at the
station and installed a Setra barometer.
Telemetry to the station has been restored. However, communication problems persist causing gaps in data acquisition.
IDA station TAU is down due to problems with communication between the digitizers and the station computer.
The long period noise at TAU ceased after day 2012,246 (Sept.2,2012).
There is excessive background noise for all channels since day 2012,173 for reasons yet to be determined.
Data quality improved and the excessive noise ceased on day 2011,227-00:00.
There is excessive background noise for all data channels since day 2011,170 for reasons yet to be determined. The noise is electromagnetic in nature and suspected from an external source. IDA engineers are working with the station operator to determine the cause.
The station is back on-line as of day 2011,145-00:39, after the station operator replaced the power supply.
The station is down as of day 2011,142-08:23 due to suspected problems with the power supply. The station operator plans to visit the station to investigate.
The station is back up as of day 2009,330 after the station operator replaced the station computer.
The station is down as of day 2009,323 due to failure of the station computer. A new computer is being prepared to ship to the station.
The station is back online as of about day 2008,269 after an IDA engineer visited the station and upgraded the data acquisition system to a Quanterra Q330HR system.
The station is down as of about day 2008,171 due to problems with the DAS that are yet to be determined. A replacement DAS is being prepared to ship to the station.
The station was down from day 062-04:26 to 063-22:42 due to problems with power to the station. The station came back online after the station operator cycled power to the ARS.
The station was down from day 352-06:25 to 354-01:41 due to a power outage caused by a faulty UPS. Data acquisition resumed after the station operator restored power to the station.
Station is back up and telemetry restored as of 12/24/04, after the station operator installed a new workstation.
Station is down due to workstation failure.
Station operator reports the vault power supply was vandalized about a week ago. The station lost power so was down for about 3/4 of a day. The station operator also replaced the DAT drive but notes there are still recording errors. Despite recording errors telemetry to the station has remained operational and data via telemetry remains complete.
Telemetry restored. NRTS data acquired as of 2003:151-02:23:23 (May 31,2003).
No NRTS data as of 2003:148-21:29:17 (May 28,2003).
Bad tape batch, replacements sent.
New EPROMs installed to correct DSP problem.
Telemetry problem resolved as of 2002133 (May 13,2002).
Station operating normally. There is no telemetry due to a disk failure on the NRTS workstation.